OpenStudy (anonymous):

number o fmiles per gallons of gas a car average varies inversely as the average speed the car travel a car gets 28 miles per gallon at 36mph how many miles per gallon will the car get at 60mph?

OpenStudy (dumbcow):

mpg = k/speed 28=k/36 solve for k k = 36*28 = 1008 mpg = 1008/speed mpg = 1008/60 = 16.8

OpenStudy (anonymous):

you multiply 28x36 than divide by 60=16.8 thanks smartcow

OpenStudy (dumbcow):

:)